The product of two positive consecutive integers is 41 more than their sum. Find the integers
Answer by checkley77(12844)

X(X+1)=X+X+1+41
X^2+X=2X+42
X^2+X-2X-42=0
X^2-X-42=0
(X-7)(X+6)
X-7=0
X=7 ANS.
X+6=0
X=-6 ANS.
PROOF:
7(7+1)=7+8+41
7*8=15+41
56=56
-6(-5)=-6-5+41
30=-11+41
30=30